# Abdulrahman Alswaidan and Jeffrey D. Varner R.F. Smith School of Chemical and Biomolecular Engineering

### 06 Mar 2026 v1

@@ Abstract
Attention heads retrieve: given a query, they return a softmax-weighted average of stored values. We show that this computation is one step of gradient descent on a classical energy function, and that Langevin sampling from the corresponding distribution yields stochastic attention: a training-free sampler controlled by a single temperature. Lowering the temperature gives exact retrieval; raising it gives open-ended generation. Because the energy gradient equals the attention map, no score network, training loop, or learned model is required. We validate on four domains (64 to 4,096 dimensions). At generation temperature, stochastic attention is 2.6 times more novel and 2.0 times more diverse than the best learned baseline (a variational autoencoder trained on the same patterns), while matching a Metropolis-corrected gold standard. A simple signal-to-noise rule selects the operating temperature for any dimension. The approach requires no architectural changes and extends naturally to retrieval-augmented generation and in-context learning.
@@ 1 Introduction
The ingredients for an answer already exist, though they have not been assembled. On one side, a line of work on associative memory, from Hopfield networks [13] through dense associative memories [16, 17] to the modern continuous formulation of Ramsauer et al. [21], has revealed that each attention head implicitly performs gradient descent on a smooth, confining energy whose minima are stored patterns. On the other side, a mature theory of Langevin dynamics [23, 30, 7] shows how to convert any such energy into a sampler for the corresponding Boltzmann distribution: one simply adds calibrated noise to the gradient update. Meanwhile, energy-based models [19, 27] and score-based diffusion methods [26, 28] have demonstrated the power of Langevin-type sampling for generation, but rely on black-box neural networks whose scores must be learned. The Energy Transformer [12] brings Hopfield energies into a deep architecture, yet remains a discriminative model that descends the energy rather than sampling from it. The classical retrieval-generation duality (Hopfield networks retrieve, Boltzmann machines sample, and both share the same energy [1]) has not been lifted to the modern continuous setting.
@@
In this paper, we close that gap. We show that applying the unadjusted Langevin algorithm to the modern Hopfield energy yields a stochastic attention update: a single iteration comprising a contraction toward the origin, a softmax attention pull toward stored memories, and an isotropic Gaussian perturbation whose magnitude is governed by the temperature. The inverse temperature $\beta$ interpolates continuously between exact retrieval ($\beta\to\infty$) and open-ended exploration ($\beta\to 0$), providing a principled generation mechanism that requires no learned score network, no training loop, and no contrastive objective. Because the gradient of the Hopfield energy is exactly the identity minus the attention map, every step of the sampler is computed by the same primitive operations as a standard attention head. In particular, $\mathbf{X}$ can be the key matrix of any pretrained attention layer, making stochastic attention a zero-shot stochastic decoding layer compatible with retrieval-augmented generation and in-context learning settings. The energy’s analytic structure (infinite differentiability, Lipschitz gradients, and a quadratic confining bound) delivers convergence guarantees that generic energy-based models cannot offer without additional assumptions. In practice, the same algorithm realizes two qualitatively distinct operating regimes: at high $\beta$ it performs structured retrieval, generating outputs that closely resemble stored patterns; once $\beta$ falls below a signal-to-noise threshold ($\mathrm{SNR}\approx 0.025$, derived in Section 4), it performs genuine generation, producing novel outputs that outperform learned baselines on both novelty and diversity. Section 2 surveys related work; Section 3 reviews the necessary background on attention, modern Hopfield networks, and Langevin dynamics; Section 4 derives the stochastic attention update, presents the algorithm, and characterizes its properties; Section 5 validates both regimes experimentally across four domains.
@@ 2 Related Work
Hopfield [13] introduced a recurrent network of $N$ binary neurons whose asynchronous dynamics minimize the quadratic energy $E(\mathbf{s})=-\frac{1}{2}\mathbf{s}^{\top}\mathbf{W}\mathbf{s}$, with $\mathbf{W}=\frac{1}{N}\sum_{i}\mathbf{m}_{i}\mathbf{m}_{i}^{\top}$ the Hebbian weight matrix. Stored patterns correspond to local minima of $E$, so retrieval proceeds by gradient descent from an initial probe. When the $K$ patterns are uncorrelated random binary vectors, Amit et al. [2] showed via a replica-symmetric mean-field analysis that reliable retrieval is possible only when the load ratio $K/N<\alpha_{c}\approx 0.138$ (where $\alpha_{c}$ denotes the critical load ratio, not to be confused with the step size $\alpha$ used later); beyond this threshold the energy landscape becomes dominated by spurious mixtures. Folli et al. [8] later gave exact finite-$N$ expressions for the bit-error probability, confirming the $\alpha_{c}$ transition and showing that the number of retrieval errors grows with $K/N$ as cross-talk noise overwhelms the coherent signal.
The modern Hopfield network program, initiated by Krotov and Hopfield [16], addresses these shortcomings by replacing the quadratic interaction with higher-order (polynomial or exponential) interaction functions. Dense associative memories with degree-$n$ polynomial energy store on the order of $N^{n-1}$ patterns, breaking the linear capacity barrier by trading pairwise for higher-order synaptic interactions, and the resulting retrieval dynamics map onto deep networks with rectified polynomial activations. At the top of this hierarchy sits the log-sum-exp (lse) energy, which achieves storage capacity that grows exponentially in the pattern dimension $N$ [21]. Ramsauer et al. [21] proved that the continuous-state retrieval map associated with the lse energy is equivalent to transformer-style attention: each attention head performs one step of gradient descent on the energy, and the fixed points of the dynamics (global, metastable, or single-pattern) correspond to different retrieval regimes. All of these works, however, focus exclusively on retrieval, that is, the convergence of iterates to stored patterns or their convex combinations. Sampling from the Boltzmann distribution defined by the same energy, which would yield a generative mechanism, is not explored. Our work fills this gap by applying Langevin dynamics to the modern Hopfield energy, converting the deterministic retrieval map into a stochastic sampler.
@@
Energy-based models (EBMs) define a probability distribution $p(\mathbf{x})\propto\exp(-E(\mathbf{x}))$ for an energy function $E$ parameterized by a neural network [19]. Training typically proceeds by contrastive divergence [10], score matching [14], or noise contrastive estimation [27]. The energy in a generic EBM is a black-box neural network, so its gradient (the score function) must be estimated or approximated. By contrast, the modern Hopfield energy has a closed-form gradient (identity minus softmax attention) requiring no auxiliary score network. Moreover, $\mathbf{X}$ is given as data rather than learned, so no training loop is needed, and the log-sum-exp structure provides analytic guarantees (smoothness, Lipschitz gradients, quadratic confinement) that generic EBMs achieve only with careful architectural design. The Energy Transformer [12] shows that Hopfield-type energies improve discriminative tasks (graph anomaly detection, image completion) but does not sample from the Boltzmann distribution; our work uses the same energy class for generation.
@@ 3 Background
The classical Hopfield network [13] stores $K$ binary patterns $\mathbf{m}_{1},\dots,\mathbf{m}_{K}\in\{-1,+1\}^{N}$ via the quadratic energy
$E_{\mathrm{classical}}(\mathbf{s})=-\tfrac{1}{2}\,\mathbf{s}^{\top}\mathbf{W}\mathbf{s}-\mathbf{b}^{\top}\mathbf{s},\qquad\mathbf{W}=\frac{1}{N}\sum_{i=1}^{K}\mathbf{m}_{i}\mathbf{m}_{i}^{\top},$ (2)
where $\mathbf{b}\in\mathbb{R}^{N}$ is an external bias (typically set to zero). Retrieval proceeds by coordinate-wise sign updates that descend $E_{\mathrm{classical}}$. The storage capacity of this network scales as $\mathcal{O}(N)$ [2], severely limiting the number of patterns.
Modern (continuous) Hopfield networks [21] overcome this bottleneck by replacing the quadratic energy with a log-sum-exp (lse) form. Let $\mathbf{X}=[\mathbf{m}_{1},\dots,\mathbf{m}_{K}]\in\mathbb{R}^{N\times K}$ be the memory matrix, $\beta>0$ an inverse temperature, and $M:=\max_{i}\|\mathbf{m}_{i}\|_{2}$. The energy is
@@
Given a smooth potential $U:\mathbb{R}^{N}\to\mathbb{R}$ with target density $p(\mathbf{x})\propto\exp(-U(\mathbf{x}))$, the (overdamped) Langevin stochastic differential equation is
@@
where $\mathbf{B}_{t}$ is standard $N$-dimensional Brownian motion. The process (6) has $p$ as its unique stationary distribution provided two regularity conditions hold [23]: (R1) $\nabla U$ is $L$-Lipschitz, i.e. $\|\nabla U(\mathbf{x})-\nabla U(\mathbf{y})\|_{2}\leq L\|\mathbf{x}-\mathbf{y}\|_{2}$; and (R2) $U$ is dissipative, i.e. $\langle\nabla U(\mathbf{x}),\mathbf{x}\rangle\geq a\|\mathbf{x}\|_{2}^{2}-b$ for constants $a>0$, $b\geq 0$. The modern Hopfield energy (3) satisfies both: it is $C^{\infty}$ with Lipschitz gradient (R1), and its quadratic lower bound (5) implies dissipativity (R2).
@@ 4.1 Stochastic Attention Update
Recall the setup from Section 3: the memory matrix $\mathbf{X}=[\mathbf{m}_{1},\dots,\mathbf{m}_{K}]\in\mathbb{R}^{N\times K}$ has $K$ stored memories as its columns, each of dimension $N$. We wish to sample from the Boltzmann distribution induced by the modern Hopfield energy (3):
@@
where $\beta>0$ is the inverse temperature and $E$ is defined in (3). To apply the ULA (7) we need $\nabla E$. The background already established the identity $\nabla E(\boldsymbol{\xi})=\boldsymbol{\xi}-\mathbf{T}(\boldsymbol{\xi})$, where $\mathbf{T}$ is the softmax attention map (4). Reparameterizing the ULA via $U=\beta E$ and rescaling time by $\beta$ (so the gradient of $E$, rather than $\beta E$, appears directly), we obtain
@@
This is the stochastic attention update. Each iteration performs three operations: a contraction toward the origin (the $(1-\alpha)\boldsymbol{\xi}_{t}$ term), a softmax-weighted pull toward stored memories (the attention term), and an isotropic Gaussian perturbation whose magnitude is governed by the temperature $1/\beta$. The complete procedure is given as pseudocode in Algorithm 1. Because the memory matrix $\mathbf{X}$ is fixed and known, each step requires only the same primitive operations as a single attention head: two matrix-vector products, one softmax, and one Gaussian draw, for a per-step cost of $\mathcal{O}(NK)$.
@@
0: Memory matrix $\mathbf{X}\in\mathbb{R}^{N\times K}$, inverse temperature $\beta>0$, step size $\alpha\in(0,1)$, number of iterations $T$, initial state $\boldsymbol{\xi}_{0}\in\mathbb{R}^{N}$
@@
3:  $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{N})$
@@ 4.2 Properties and Limiting Behavior
Let $\sigma_{\max}=\|\mathbf{X}\|_{\mathrm{op}}$ denote the largest singular value of the memory matrix. The modern Hopfield energy (3) has Lipschitz-continuous gradient with constant $L=1+\beta\sigma_{\max}^{2}/2$ and satisfies the dissipativity condition $\langle\nabla E(\boldsymbol{\xi}),\boldsymbol{\xi}\rangle\geq\|\boldsymbol{\xi}\|_{2}^{2}/2-M^{2}/2$ for all $\boldsymbol{\xi}\in\mathbb{R}^{N}$.
The Lipschitz constant grows linearly in $\beta$: the energy landscape sharpens at low temperature and step-size selection becomes more delicate. The Hessian satisfies $\nabla^{2}E\succeq(1-\beta\sigma_{\max}^{2}/2)\mathbf{I}$, so $E$ is strictly convex when $\beta\sigma_{\max}^{2}<2$, yielding the following convergence guarantee.
@@
When $\beta\sigma_{\max}^{2}<2$, the potential $U=\beta E$ is $m$-strongly convex with $m=\beta(1-\beta\sigma_{\max}^{2}/2)$ and has $L_{U}$-Lipschitz gradient with $L_{U}=\beta(1+\beta\sigma_{\max}^{2}/2)$. The iterates of Algorithm 1 then converge to $p_{\beta}$ in $W_{2}$ at a geometric rate, with a discretization bias that vanishes as $\alpha\to 0$ [4, 7]. For arbitrary $\beta>0$ the continuous-time Langevin diffusion is geometrically ergodic [23]; the ULA discretization approximates $p_{\beta}$ with bias $O(\alpha)$.
Scope of the convergence guarantee.
Corollary 2’s geometric-rate guarantee requires $\beta\sigma_{\max}^{2}<2$, i.e., a log-concave energy. In our experiments $\sigma_{\max}^{2}\approx 2$33, so this condition holds only for $\beta\lesssim 1$, far below the operating range. At $\beta\geq 200$ the energy is multimodal and the log-concave rate bound no longer applies; deriving explicit mixing-time bounds in this regime is an open problem. Two complementary guarantees nonetheless support the operating regime: the continuous-time Langevin SDE is geometrically ergodic for any $\beta>0$ by Proposition 1 [23], and the MALA acceptance rate of 99.2% at $\alpha{=}0.01$ (Table 1) confirms negligible ULA discretization bias in practice.
@@
Because $\mathbf{X}$ is given as data, there are no learned parameters: no training loop, no contrastive divergence, and no score-matching objective. The confining bound (5) keeps iterates concentrated near $\mathrm{conv}\{\mathbf{m}_{1},\dots,\mathbf{m}_{K}\}$ with excursions controlled by $\sqrt{2\alpha/\beta}$; Section 5 derives a dimension-independent $\beta$-selection rule from the per-step signal-to-noise ratio.

### 10 Mar 2026 v2

@@ Abstract
Attention heads retrieve: given a query, they return a softmax-weighted average of stored values. We show that this computation is one step of gradient descent on a classical energy function, and that Langevin sampling from the corresponding distribution yields stochastic attention: a training-free sampler controlled by a single temperature. Lowering the temperature gives exact retrieval; raising it gives open-ended generation. Because the energy gradient equals the attention map, no score network, training loop, or learned model is required. We derive a closed-form entropy inflection condition that identifies the retrieval-to-generation transition temperature for any memory geometry, with a scaling law $\beta^{*}\!\sim\!\sqrt{d}$ for random patterns. We validate on five domains (64 to 4,096 dimensions). On MNIST digit images, stochastic attention is $2.6{\times}$ more novel and $2.0{\times}$ more diverse than the best learned baseline (a VAE trained on the same patterns), while matching a Metropolis-corrected gold standard. On protein sequences from the Pfam RRM family, the generation regime achieves $6.9{\times}$ lower amino acid composition divergence than the VAE (KL =0.060=0.060 vs. $0.416$) at matched novelty, demonstrating that the training-free score function preserves family-level fidelity that learned models lose. A denoising diffusion baseline (DDPM) fails across all memory sizes tested ($K=100$ to 3,5003{,}500), producing samples indistinguishable from isotropic noise. The approach requires no architectural changes to the underlying attention mechanism.
@@ 1 Introduction
The ingredients for an answer already exist, though they have not been assembled. On one side, a line of work on associative memory, from Hopfield networks [15] through dense associative memories [19, 20] to the modern continuous formulation of Ramsauer et al. [26], has revealed that each attention head implicitly performs gradient descent on a smooth, confining energy whose minima are stored patterns. On the other side, a mature theory of Langevin dynamics [28, 35, 8] shows how to convert any such energy into a sampler for the corresponding Boltzmann distribution: one simply adds calibrated noise to the gradient update. Meanwhile, energy-based models [22, 32] and score-based diffusion methods [31, 33] have demonstrated the power of Langevin-type sampling for generation, but rely on black-box neural networks whose scores must be learned. The Energy Transformer [14] brings Hopfield energies into a deep architecture, yet remains a discriminative model that descends the energy rather than sampling from it. The classical retrieval-generation duality (Hopfield networks retrieve, Boltzmann machines sample, and both are defined by the same quadratic interaction energy over binary states [1]) has not been lifted to the modern continuous setting.
@@
In this paper, we close that gap. We show that applying the unadjusted Langevin algorithm to the modern Hopfield energy yields a stochastic attention update: a single iteration comprising a contraction toward the origin, a softmax attention pull toward stored memories, and an isotropic Gaussian perturbation whose magnitude is governed by the temperature. The inverse temperature $\beta$ interpolates continuously between exact retrieval ($\beta\to\infty$) and open-ended exploration ($\beta\to 0$), providing a principled generation mechanism that requires no learned score network, no training loop, and no contrastive objective. Because the gradient of the Hopfield energy is exactly the identity minus the attention map, every step of the sampler is computed by the same primitive operations as a standard attention head. In particular, $\mathbf{X}$ can be the key matrix of any pretrained attention layer, making stochastic attention a zero-shot stochastic decoding layer. When the attention head uses learned $Q$, $K$, $V$ projections operating on hidden states rather than raw patterns, the connection requires projecting through these learned transformations; we leave this extension, along with applications to retrieval-augmented generation and in-context learning, to future work. The energy’s analytic structure (infinite differentiability, Lipschitz gradients, and a quadratic confining bound) delivers convergence guarantees that generic energy-based models cannot offer without additional assumptions. In practice, the same algorithm realizes two qualitatively distinct operating regimes: at high $\beta$ it performs structured retrieval, generating outputs that closely resemble stored patterns; once $\beta$ falls below a critical signal-to-noise ratio $\mathrm{SNR}^{*}$ whose scaling is derived from the entropy inflection of the attention distribution (Section 4), it performs genuine generation, producing novel outputs that outperform learned baselines on both novelty and diversity while preserving domain-level fidelity (e.g., $6.9{\times}$ lower amino acid composition divergence than a VAE on protein sequences). The individual ingredients, the Hopfield energy gradient [26] and the Langevin algorithm [28], are known. The contribution lies in their synthesis and in the analysis it enables, which is inaccessible from either ingredient alone. First, because the Hopfield energy gradient is exactly the identity minus the attention map, the score function is available in closed form at the cost of a single attention evaluation; this eliminates the score matching or contrastive divergence required by standard energy-based and score-based generative models, and converts any pretrained attention head into a sampler with no additional training. Second, Ramsauer et al. studied the energy only for retrieval ($\beta\to\infty$) and the Langevin literature does not know about attention; by analyzing the energy as a sampling target we characterize the full temperature spectrum, and the entropy inflection analysis (Proposition 3) identifies the critical $\beta^{*}$ at which the sampler transitions from retrieval to generation, with a scaling law $\beta^{*}\sim\sqrt{d}$ that yields a dimension-dependent SNR criterion for temperature selection. Third, the smoothness, dissipativity, and condition number of the Hopfield energy as a function of $\beta$ and $K/d$ had not been characterized for the purpose of sampling guarantees; we provide these bounds (Proposition 1) and show where they apply and where they break down. Section 2 surveys related work; Section 3 reviews the necessary background on attention, modern Hopfield networks, and Langevin dynamics; Section 4 derives the stochastic attention update, presents the algorithm, and characterizes its properties; Section 5 validates both regimes experimentally across five domains.
@@ 2 Related Work
Hopfield [15] introduced a recurrent network of $d$ binary neurons whose asynchronous dynamics minimize the quadratic energy $E(\mathbf{s})=-\frac{1}{2}\mathbf{s}^{\top}\mathbf{W}\mathbf{s}$, with $\mathbf{W}=\frac{1}{d}\sum_{i}\mathbf{m}_{i}\mathbf{m}_{i}^{\top}$ the Hebbian weight matrix. Stored patterns correspond to local minima of $E$, so retrieval proceeds by gradient descent from an initial probe. When the $K$ patterns are uncorrelated random binary vectors, Amit et al. [2] showed via a replica-symmetric mean-field analysis that reliable retrieval is possible only when the load ratio $K/d<\alpha_{c}\approx 0.138$ (where $\alpha_{c}$ denotes the critical load ratio, not to be confused with the step size $\alpha$ used later); beyond this threshold the energy landscape becomes dominated by spurious mixtures. Folli et al. [9] later gave exact finite-$d$ expressions for the bit-error probability, confirming the $\alpha_{c}$ transition and showing that the number of retrieval errors grows with $K/d$ as cross-talk noise overwhelms the coherent signal.
The modern Hopfield network program, initiated by Krotov and Hopfield [19], addresses these shortcomings by replacing the quadratic interaction with higher-order (polynomial or exponential) interaction functions. Dense associative memories with degree-$n$ polynomial energy store on the order of $d^{n-1}$ patterns, breaking the linear capacity barrier by trading pairwise for higher-order synaptic interactions, and the resulting retrieval dynamics map onto deep networks with rectified polynomial activations. At the top of this hierarchy sits the log-sum-exp (lse) energy, which achieves storage capacity that grows exponentially in the pattern dimension $d$ [26]. Ramsauer et al. [26] proved that the continuous-state retrieval map associated with the lse energy is equivalent to transformer-style attention: each attention head performs one step of gradient descent on the energy, and the fixed points of the dynamics (global, metastable, or single-pattern) correspond to different retrieval regimes. All of these works, however, focus exclusively on retrieval, that is, the convergence of iterates to stored patterns or their convex combinations. Sampling from the Boltzmann distribution defined by the same energy, which would yield a generative mechanism, is not explored. Our work fills this gap by applying Langevin dynamics to the modern Hopfield energy, converting the deterministic retrieval map into a stochastic sampler.
@@
Energy-based models (EBMs) define a probability distribution $p(\mathbf{x})\propto\exp(-E(\mathbf{x}))$ for an energy function $E$ parameterized by a neural network [22]. Training typically proceeds by contrastive divergence [12], score matching [16], or noise contrastive estimation [10, 32]. The energy in a generic EBM is a black-box neural network, so its gradient (the score function) must be estimated or approximated. By contrast, the modern Hopfield energy has a closed-form gradient (identity minus softmax attention) requiring no auxiliary score network. Moreover, $\mathbf{X}$ is given as data rather than learned, so no training loop is needed, and the log-sum-exp structure provides analytic guarantees (smoothness, Lipschitz gradients, quadratic confinement) that generic EBMs achieve only with careful architectural design. The Energy Transformer [14] shows that Hopfield-type energies improve discriminative tasks (graph anomaly detection, image completion) but does not sample from the Boltzmann distribution; our work uses the same energy class for generation.
@@ 3 Background
The classical Hopfield network [15] stores $K$ binary patterns $\mathbf{m}_{1},\dots,\mathbf{m}_{K}\in\{-1,+1\}^{d}$ via the quadratic energy
$E_{\mathrm{classical}}(\mathbf{s})=-\tfrac{1}{2}\,\mathbf{s}^{\top}\mathbf{W}\mathbf{s}-\mathbf{b}^{\top}\mathbf{s},\qquad\mathbf{W}=\frac{1}{d}\sum_{i=1}^{K}\mathbf{m}_{i}\mathbf{m}_{i}^{\top},$ (2)
where $\mathbf{b}\in\mathbb{R}^{d}$ is an external bias (typically set to zero). Retrieval proceeds by coordinate-wise sign updates that descend $E_{\mathrm{classical}}$. The storage capacity of this network scales as $\mathcal{O}(d)$ [2], severely limiting the number of patterns.
Modern (continuous) Hopfield networks [26] overcome this bottleneck by replacing the quadratic energy with a log-sum-exp (lse) form. Let $\mathbf{X}=[\mathbf{m}_{1},\dots,\mathbf{m}_{K}]\in\mathbb{R}^{d\times K}$ be the memory matrix, $\beta>0$ an inverse temperature, and $M:=\max_{i}\|\mathbf{m}_{i}\|_{2}$. The energy is
@@
Given a smooth potential $U:\mathbb{R}^{d}\to\mathbb{R}$ with target density $p(\mathbf{x})\propto\exp(-U(\mathbf{x}))$, the (overdamped) Langevin stochastic differential equation is
@@
where $\mathbf{B}_{t}$ is standard $d$-dimensional Brownian motion. The process (6) has $p$ as its unique stationary distribution provided two regularity conditions hold [28]: (R1) $\nabla U$ is $L$-Lipschitz, i.e. $\|\nabla U(\mathbf{x})-\nabla U(\mathbf{y})\|_{2}\leq L\|\mathbf{x}-\mathbf{y}\|_{2}$; and (R2) $U$ is dissipative, i.e. $\langle\nabla U(\mathbf{x}),\mathbf{x}\rangle\geq a\|\mathbf{x}\|_{2}^{2}-b$ for constants $a>0$, $b\geq 0$. The modern Hopfield energy (3) satisfies both: it is $C^{\infty}$ with Lipschitz gradient (R1), and its quadratic lower bound (5) implies dissipativity (R2).
@@ 4.1 Stochastic Attention Update
Recall the setup from Section 3: the memory matrix $\mathbf{X}=[\mathbf{m}_{1},\dots,\mathbf{m}_{K}]\in\mathbb{R}^{d\times K}$ has $K$ stored memories as its columns, each of dimension $d$. We wish to sample from the Boltzmann distribution induced by the modern Hopfield energy (3):
@@
where $\beta>0$ is the inverse temperature and $E$ is defined in (3). To apply the ULA (7) we need $\nabla E$. The background already established the identity $\nabla E(\boldsymbol{\xi})=\boldsymbol{\xi}-\mathbf{T}(\boldsymbol{\xi})$, where $\mathbf{T}$ is the softmax attention map (4). Applying the ULA to the potential $U=\beta E$ with step size $\tilde{\alpha}$ gives $\boldsymbol{\xi}_{t+1}=\boldsymbol{\xi}_{t}-\tilde{\alpha}\beta\nabla E+\sqrt{2\tilde{\alpha}}\,\boldsymbol{\epsilon}_{t}$. Substituting $\tilde{\alpha}=\alpha/\beta$ (so the effective step for $\nabla E$ is $\alpha$ and the noise scale becomes $\sqrt{2\alpha/\beta}$), we obtain
@@
This is the stochastic attention update. Each iteration performs three operations: a contraction toward the origin (the $(1-\alpha)\boldsymbol{\xi}_{t}$ term), a softmax-weighted pull toward stored memories (the attention term), and an isotropic Gaussian perturbation whose magnitude is governed by the temperature $1/\beta$. The complete procedure is given as pseudocode in Algorithm 1. Because the memory matrix $\mathbf{X}$ is fixed and known, each step requires only the same primitive operations as a single attention head: two matrix-vector products, one softmax, and one Gaussian draw, for a per-step cost of $\mathcal{O}(dK)$.
@@
0: Memory matrix $\mathbf{X}\in\mathbb{R}^{d\times K}$, inverse temperature $\beta>0$, step size $\alpha\in(0,1)$, number of iterations $T$, initial state $\boldsymbol{\xi}_{0}\in\mathbb{R}^{d}$
@@
3:  $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$
@@ 4.2 Properties and Limiting Behavior
Let $\sigma_{\max}=\|\mathbf{X}\|_{\mathrm{op}}$ denote the largest singular value of the memory matrix. The modern Hopfield energy (3) has Lipschitz-continuous gradient with constant $L=1+\beta\sigma_{\max}^{2}/2$ and satisfies the dissipativity condition $\langle\nabla E(\boldsymbol{\xi}),\boldsymbol{\xi}\rangle\geq\|\boldsymbol{\xi}\|_{2}^{2}/2-M^{2}/2$ for all $\boldsymbol{\xi}\in\mathbb{R}^{d}$.
The Lipschitz constant grows linearly in $\beta$: the energy landscape sharpens at low temperature and step-size selection becomes more delicate. The Hessian satisfies $\nabla^{2}E\succeq(1-\beta\sigma_{\max}^{2}/2)\mathbf{I}$, so $E$ is strictly convex when $\beta\sigma_{\max}^{2}<2$. This condition characterizes the convex regime of the Hopfield energy and yields a standard ULA convergence result:
@@
When $\beta\sigma_{\max}^{2}<2$, the potential $U=\beta E$ is $m$-strongly convex with $m=\beta(1-\beta\sigma_{\max}^{2}/2)$ and has $L_{U}$-Lipschitz gradient with $L_{U}=\beta(1+\beta\sigma_{\max}^{2}/2)$. The iterates of Algorithm 1 then converge to $p_{\beta}$ in $W_{2}$ at a geometric rate, with a discretization bias that vanishes as $\alpha\to 0$ [5, 8].
Beyond the convex regime.
Corollary 2 applies when $\beta\sigma_{\max}^{2}<2$. In our experiments $\sigma_{\max}^{2}\approx 2$33, so this condition holds only for $\beta\lesssim 1$, well below the operating range ($\beta\in\{200,2000\}$). We state this boundary explicitly because it delineates where the Hopfield energy transitions from unimodal (one global basin) to multimodal (one local minimum per stored pattern): the condition $\beta\sigma_{\max}^{2}=2$ is the onset of non-convexity, beyond which energy barriers separate stored memories.
Proving mixing-time bounds in the multimodal regime ($\beta\sigma_{\max}^{2}>2$) is a fundamental open problem, not specific to our setting. For general non-convex energies, the mixing time of Langevin dynamics scales exponentially in the barrier height divided by the temperature [3]. In the Hopfield energy at high $\beta$, the barriers between stored patterns are $O(\beta)$ while the temperature is $1/\beta$, so inter-basin mixing is exponentially slow. This is consistent with our empirical observation that a single chain at $\beta{=}2000$ does not cross between basins (single-chain diversity $0.282$; Appendix F).
However, the practical situation is more favorable than the worst-case theory suggests. Within a single basin the energy is locally strongly convex (the Hessian near each minimum is positive definite), so intra-basin mixing is fast. The multi-chain protocol (Section 5) is designed to exploit this: by initializing 30 chains near distinct stored patterns, we obtain coverage across basins through initialization and rely on fast local mixing within each basin. At $\beta{=}200$, the barriers are low enough that single chains do cross basins (single-chain diversity $0.796$, exceeding the multi-chain $\beta{=}2000$ value of $0.600$).
Two guarantees hold for all $\beta>0$ regardless of convexity. First, the continuous-time Langevin SDE is geometrically ergodic by the dissipativity of Proposition 1 [28], ensuring convergence to the correct Boltzmann target (though without an explicit rate). Second, the MALA acceptance rate of $99.2\%$ at $\alpha{=}0.01$ (Table 1) confirms that ULA discretization bias is negligible in practice, and the step-size sweep (Appendix K) maps where this approximation breaks down.
@@
Because $\mathbf{X}$ is given as data, there are no learned parameters: no training loop, no contrastive divergence, and no score-matching objective. The confining bound (5) keeps iterates concentrated near $\mathrm{conv}\{\mathbf{m}_{1},\dots,\mathbf{m}_{K}\}$ with excursions controlled by $\sqrt{2\alpha/\beta}$; Section 4.3 derives a $\beta$-selection rule from the entropy of the attention distribution.
4.3 Phase Transition and Temperature Selection
The attention entropy $H(\beta)=-\sum_{k}p_{k}\log p_{k}$, where $p_{k}=\operatorname{softmax}(\beta\mathbf{X}^{\top}\boldsymbol{\xi})_{k}$, quantifies the selectivity of the attention distribution for a given state $\boldsymbol{\xi}$. At $\beta=0$ the distribution is uniform and $H=\log K$; as $\beta\to\infty$ it concentrates on the nearest memory and $H\to 0$. The following identity governs the rate of this transition.
Proposition 3.
For fixed $\boldsymbol{\xi}\in\mathbb{R}^{d}$, let $e_{k}=\mathbf{m}_{k}^{\top}\boldsymbol{\xi}$ and $\mathbf{p}=\operatorname{softmax}(\beta\mathbf{e})$. Then
$\frac{dH}{d\beta}\;=\;-\beta\,\mathrm{Var}_{\mathbf{p}}(e),$ (11)
where $\mathrm{Var}_{\mathbf{p}}(e)=\sum_{k}p_{k}\,e_{k}^{2}-\bigl(\sum_{k}p_{k}\,e_{k}\bigr)^{\!2}$.
Proof.
Write $\log p_{k}=\beta e_{k}-\log Z$ with $Z=\sum_{j}\exp(\beta e_{j})$, so that $H=-\beta\langle e\rangle_{\mathbf{p}}+\log Z$. Differentiating and using the standard identities $d(\log Z)/d\beta=\langle e\rangle_{\mathbf{p}}$ and $d\langle e\rangle_{\mathbf{p}}/d\beta=\mathrm{Var}_{\mathbf{p}}(e)$ yields the result. ∎
Since $\mathrm{Var}_{\mathbf{p}}(e)\geq 0$, entropy decreases monotonically with $\beta$. At $\beta=0$ the derivative vanishes because the prefactor $\beta$ is zero; as $\beta\to\infty$ it again vanishes because the variance collapses. The maximum rate of entropy loss therefore occurs at an intermediate inflection point $\beta^{*}$ satisfying $d^{2}H/d\beta^{2}=0$:
$\mathrm{Var}_{\mathbf{p}}(e)\big|_{\beta^{*}}\;=\;-\beta^{*}\;\mu_{3}\big|_{\beta^{*}},$ (12)
where $\mu_{3}=\langle(e-\langle e\rangle_{\mathbf{p}})^{3}\rangle_{\mathbf{p}}$ is the third central moment of the similarities under the attention distribution. This inflection marks the retrieval–generation phase boundary: below $\beta^{*}$, attention remains diffuse (generation); above it, attention concentrates on individual memories (retrieval).
Scaling for random memories.
When the $K$ memories are drawn uniformly on the unit sphere $\mathbb{S}^{d-1}$, the similarities $e_{k}$ are approximately iid with $\mathrm{Var}(e_{k})=1/d$. The softmax logits $\beta e_{k}$ have standard deviation $\beta/\sqrt{d}$; the transition from near-uniform to concentrated attention occurs when this standard deviation reaches order unity, giving
$\beta^{*}\;\sim\;\sqrt{d}.$ (13)
Substituting into the per-step signal-to-noise ratio (Eq. 15) yields a scaling law for the critical SNR:
$\mathrm{SNR}^{*}\;=\;\sqrt{\frac{\alpha\,\beta^{*}}{2d}}\;\sim\;\sqrt{\frac{\alpha}{2\sqrt{d}}}\,.$ (14)
At d=64d{=}64 and $\alpha{=}0.01$ this evaluates to $\mathrm{SNR}^{*}=\sqrt{0.01/16}=0.025$, exactly matching the empirical transition observed in Section 5. The threshold is not a universal constant: it scales as $d^{-1/4}$ and depends on the pattern geometry through $\mathrm{Var}(e_{k})$. For structured data the effective variance may differ substantially from $1/d$, but Eq. (12) provides a fully data-dependent criterion that applies in any domain: evaluate $d^{2}H/d\beta^{2}$ over a sweep of $\beta$ values and locate the zero crossing.